OpenStudy (anonymous):

how many electrons would it take to equal the mass of a proton using two significant figures

OpenStudy (jfraser):

You're looking for a ratio that looks like:\[\frac {\# of electrons}{1 photon}\] If you know the mass of each electron, and the mass of each photon, you can do the same thing with the masses:\[\frac {\frac{1.6726*10^-{24}g}{1 proton}}{\frac {9.109*10^{-28}g}{1electron}}\] I get about 1839 electrons per photon mass. To 2 sigfigs, that's 1800.